From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sun Mar 16 2003 - 16:22:04 EST
------------------------------------------------------------ Date: Sun Mar 16 16:22:04 2003, Shift: 08:00-16:00 Supervisor: sjs, with: rl Log book#14 pages 65 - 65 Shift Summary: ------------------------------------------------------------ Finished run 8420 at 8:46 when we voted with Phenex to dump beam (zdc rate < 1/3 initial rate). Various rhic machine problems resulted in no new stores during shift. It is a beutiful day at BNL so, of course, the air conditioner for the FEH tripped, taking with it the clean power. After several hopeful "repairs", the service personnel inform us that the fan is bad and will have to be replaced. It is not yet known whether there is a replacement fan at BNL. We have been supplied with a large floor fan and are trying to cool the electronics with it. Also, the support people have rigged the air circulation system to run without AC, pushing some air through the trailer. Hopefully it will turn cold overnight...So far, the trailer is still warm. ------------------------------------------------------------
This archive was generated by hypermail 2.1.5 : Sun Mar 16 2003 - 16:26:36 EST